OpenStudy (anonymous):

What is the smallest solution of the equation x^4 - 34x^2 + 225 = 0?

OpenStudy (mertsj):

(x^2-25)(x^2-9)=0 x^2=25 or x^2=x = 5, -5, 3, -3 Smallest solution is -5

OpenStudy (anonymous):

What's your first step

OpenStudy (mertsj):

To factor the trinomial you posted.

OpenStudy (anonymous):

How do you factor it?

OpenStudy (mertsj):

I don't understand you question. Don't you see the factors I posted? That's how you factor it.

OpenStudy (anonymous):

No, that's the result you get after factoring it, its factors. Show me how you came up with the two binomials

OpenStudy (anonymous):

not sure

OpenStudy (mertsj):

Do you know how to factor? Factor this: \[x^2-34x+225\]

OpenStudy (mertsj):

Try to think of two numbers. Their product is 225 and their sum is -34

OpenStudy (anonymous):

yes i know how to factor quadratic equations

OpenStudy (mertsj):

Then factor the example I gave you and tell me what the factors are.

OpenStudy (anonymous):

(x-25)(x-9)

OpenStudy (mertsj):